A coin is tossed and a 6-sided die is rolled. What is the probability that the coin shows head and the dies shows a 4?
Answer by jim_thompson5910(35256)

P(heads AND rolling a 4) = P(heads) * P(rolling a 4)



P(heads AND rolling a 4) = (1/2) * (1/6)


P(heads AND rolling a 4) = (1*1)/(2*6)


P(heads AND rolling a 4) = 1/12